# Landau.bib

### v1

@@ 3.2.2 Approximated chain rule
$(\partial_{t}+v\cdot\nabla_{x})\mu^{\alpha,\beta,\delta,\theta}_{t}+\mathsf{E}^{\alpha,\beta,\delta,\theta}_{t}+\frac{1}{2}(\tilde{\nabla}\cdot\mathcal{U}^{\alpha,\delta,\theta})^{\beta}=0,$ (3.30)
where (~$(\tilde{\nabla}\cdot\mathcal{U}^{\alpha,\delta})^{\beta}=(\tilde{\nabla}\cdot\mathcal{U}^{\alpha,\delta}_{t})*_{(v,v_{*})}M_{\beta}$. The error term $\mathsf{E}^{\alpha,\beta,\delta,\theta}_{t}$ is given by
$\mathsf{E}^{\alpha,\beta,\delta}_{t}=(1-\theta)\int_{\mathbb{R}^{3}}M_{\beta}(u)u\cdot\nabla_{x}f^{\alpha,\delta}_{t}(x,v-u){\,\rm d}u.$
Indeed, we verify (3.30) in the weak formulation with test function $\phi\in C^{\infty}_{c}(\mathbb{R}^{3}\times\mathbb{R}^{3})$:
$\frac{d}{dt}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi{\,\rm d}\mu_{t}^{\alpha,\beta,\delta,\theta}=(1-\theta)\frac{d}{dt}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi^{\alpha,\beta}{\,\rm d}\mu^{\delta}_{t}+\theta\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi\partial_{t}g{\,\rm d}x{\,\rm d}v,$
@@
$-\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}v\cdot\nabla_{x}\phi{\,\rm d}\mu^{\alpha,\beta,\delta,\theta}_{t}$
[…]
@@
[…]
$\quad+(1-\theta)\int_{\mathbb{R}^{3\times 3}}\phi(v)M_{\beta}(u)u\cdot\nabla_{x}f^{\alpha,\delta}_{t}(x,v-u){\,\rm d}u{\,\rm d}x{\,\rm d}v,$
@@
$\int_{\mathbb{R}^{12}}\tilde{\nabla}\phi^{\alpha,\beta}{\,\rm d}\mathcal{U}^{\delta}_{t}$ $=\int_{\mathbb{R}^{12}}\tilde{\nabla}\phi^{\beta}{\,\rm d}\mathcal{U}^{\alpha,\delta}_{t}=-\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi(\tilde{\nabla}\cdot\mathcal{U}^{\alpha,\delta}_{t})*_{(v,v_{*})}M_{\beta},$
where we use $[\Pi_{(v-v_{*})^{\perp}},M_{\alpha}*_{(x,x_{*})}]=0$ and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes commutator.
@@
[…]
where the bounds on $|\log f^{\alpha,\beta,\delta,\theta}|$ are given in (3.16). The right-hand side of the above inequality is integrable on $\mathbb{R}^{3}\times\mathbb{R}^{3}$ by (3.13). We use the dominated convergence theorem to exchange the time derivative and the integral over $\mathbb{R}^{3}\times\mathbb{R}^{3}$ to derive
$\frac{d}{dr}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f^{\alpha,\beta,\delta,\theta}_{r}\log f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v$
[…]
@@
$\frac{d}{dt}\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{t})$ $=-\int_{\mathbb{R}^{6}}\mathsf{E}^{\alpha,\beta,\delta,\theta}_{t}\log f^{\alpha,\beta,\delta,\theta}_{t}{\,\rm d}x{\,\rm d}v$ (3.31)
$\quad+\frac{1}{2}\int_{\mathbb{R}^{12}}U^{\alpha,\delta,\theta}_{t}\cdot\tilde{\nabla}(\log f^{\alpha,\beta,\delta,\theta}_{t})^{\beta}{\,\rm d}\eta.$
since $\int_{\mathbb{R}^{6}}v\cdot\nabla_{x}f^{\alpha,\beta,\delta,\theta}_{t}\log f^{\alpha,\beta,\delta,\theta}_{t}{\,\rm d}x{\,\rm d}v=0$.
@@
$\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{t})-\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{s})=$ $-\int_{s}^{t}\int_{\mathbb{R}^{6}}\mathsf{E}^{\alpha,\beta,\delta,\theta}_{r}\log f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v{\,\rm d}r$ (3.32)
$+\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}^{12}}U^{\alpha,\delta,\theta}_{r}\cdot\tilde{\nabla}(\log f^{\alpha,\beta,\delta,\theta}_{r})^{\beta}{\,\rm d}\eta{\,\rm d}r$
@@
$f^{\alpha,\beta,\delta,\theta}_{t}\to f^{\alpha,\beta,\theta}_{t}\quad\text{uniformly on }(0,T)\times\mathbb{R}^{3}\times\mathbb{R}^{3}.$
@@
where $\Omega\subset\mathbb{R}^{3}\times\mathbb{R}^{3}$ bounded.
@@
[…]
@@
[…]
@@
$|\nabla_{v}\log f^{\alpha,\beta,\delta}_{t}|\lesssim\beta^{-1}\quad\text{and}\quad|\tilde{\nabla}(\log f^{\alpha,\beta,\delta}_{t})^{\beta}|\lesssim_{\beta}|v-v_{*}|^{1+\frac{\gamma}{2}}.$ (3.33)
@@
|~$|\tilde{\nabla}(\log f^{\alpha,\beta,\delta}_{t})^{\beta}|\lesssim_{\beta,\gamma}|v|^{1+\frac{\gamma}{2}}+|v_{*}|^{1+\frac{\gamma}{2}}.$
@@
The proof follows [carrillo2024landau, Lemma 31 & 32]. We note that
@@
$|\nabla_{v}\log f^{\alpha,\beta,\delta}_{t}|$ $=\Big{|}\frac{\int_{\mathbb{R}^{3}}f^{\alpha,\delta}(x,w)\nabla_{v}M_{\beta}(v-w){\,\rm d}w}{f^{\alpha,\beta,\delta}_{t}}\Big{|}$
$\lesssim\beta^{-1}\Big{|}\frac{\int_{\mathbb{R}^{3}}f^{\alpha,\delta}(x,w)M_{\beta}(v-w){\,\rm d}w}{f^{\alpha,\beta,\delta}_{t}}\Big{|}\lesssim\beta^{-1}.$
@@
~$\tilde{\nabla}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}=$ $|v-v_{*}|^{1+\frac{\gamma}{2}}\Pi_{(v-v_{*})^{\perp}}\big{(}\nabla_{v}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}-\nabla_{v_{*}}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}\big{)}.$
@@
    $\lim_{\theta\to 0}\lim_{\delta\to 0}\int_{s}^{t}\int_{\mathbb{R}^{12}}\tilde{\nabla}(\log f_{r}^{\alpha,\beta,\delta,\theta})^{\beta}\cdot U_{r}^{\alpha,\delta,\theta}{\,\rm d}\eta{\,\rm d}r=\int_{s}^{t}\int_{\mathbb{R}^{12}}\tilde{\nabla}(\log f_{r}^{\alpha,\beta})^{\beta}\cdot U_{r}^{\alpha}{\,\rm d}\eta{\,\rm d}r.$
@@
    $(1-\theta)\Big{|}\int_{s}^{t}\int_{\mathbb{R}^{12}}\tilde{\nabla}(\log f_{r}^{\alpha,\beta,\theta})^{\beta}\cdot\Big{(}U_{r}^{\alpha,\delta}-U_{r}^{\alpha}\Big{)}{\,\rm d}\eta{\,\rm d}r\Big{|}\to 0\quad\text{and}$ (3.35)
    $(1-\theta)\Big{|}\int_{s}^{t}\int_{\mathbb{R}^{12}}\Big{(}\tilde{\nabla}(\log f_{r}^{\alpha,\beta,\delta,\theta})^{\beta}-\tilde{\nabla}(\log f_{r}^{\alpha,\beta,\theta})^{\beta}\Big{)}\cdot U_{r}^{\alpha,\delta}{\,\rm d}\eta{\,\rm d}r\Big{|}\to 0.$ (3.36)
@@
    |~$|\tilde{\nabla}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}|\lesssim_{\beta}|v-v_{*}|^{1+\frac{\gamma}{2}},\quad|U^{\alpha,\beta,\delta}||v-v_{*}|^{1+\frac{\gamma}{2}}\in L^{1}([0,T]\times\mathbb{R}^{12}).$ (3.37)
@@
    We define the ball $B^{d}_{k}(\cdot)\subset\mathbb{R}^{3d}$ as the ball centred at $\cdot\in\mathbb{R}^{3d}$ with radius $k$.
@@
    |~$|\tilde{\nabla}(\log f_{r}^{\alpha,\beta,\delta,\theta})^{\beta}-\tilde{\nabla}(\log f_{r}^{\alpha,\beta,\theta})^{\beta}|$
@@
[…]
@@
    ×$\quad\times\||v-v_{*}|^{1+\frac{\gamma}{2}}U^{\alpha,\delta}\|_{L^{1}([0,T]\times\mathbb{R}^{12})}\to 0\quad\text{as}\quad\delta\to 0,$
@@
    We use the dominated convergence Theorem 3.12 to show the $\theta$-limit. Notice that the bounds (3.37) imply that |~$\big{|}\tilde{\nabla}\big{(}\log f^{\alpha,\beta,\theta}\big{)}^{\beta}\cdot U^{\alpha,\theta}\big{|}\lesssim\beta^{-1}|U^{\alpha}|$.
@@
    $\lim_{\beta\to 0}\int_{s}^{t}\int_{\mathbb{R}^{12}}\tilde{\nabla}(\log f^{\alpha,\beta})^{\beta}\cdot U^{\alpha}{\,\rm d}\eta{\,\rm d}r=\int_{s}^{t}\int_{\mathbb{R}^{12}}\tilde{\nabla}(\log f^{\alpha})\cdot U^{\alpha}{\,\rm d}\eta{\,\rm d}r.$ (3.38)
    We follow [carrillo2024landau]. By Cauchy–Schwarz inequality, we have
    $\Big{|}\int_{s}^{t}\int_{\mathbb{R}^{12}}\tilde{\nabla}(\log f^{\alpha,\beta})^{\beta}\cdot U^{\alpha}{\,\rm d}\eta{\,\rm d}r\Big{|}$
    $\leq\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{12}}\kappa f^{\alpha}f^{\alpha}_{*}|\tilde{\nabla}(\log f^{\alpha,\beta})^{\beta}|^{2}{\,\rm d}\eta{\,\rm d}t+\frac{1}{2}\int_{0}^{T}\mathcal{A}(f^{\alpha}){\,\rm d}t,$
@@
    Similar to (2.16), the three-dimensional cross-product identity (2.5) implies that
    |~$\Big{|}\tilde{\nabla}(\log f^{\alpha,\beta})^{\beta}\Big{|}^{2}$
    $=|v-v_{*}|^{\gamma}\Big{|}(v-v_{*})\times(\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}-\nabla_{v_{*}}\log f^{\alpha,\beta}_{*}*_{v_{*}}M_{\beta})\Big{|}^{2}.$
@@
    $\Big{|}(v-v_{*})\times(\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}-\nabla_{v_{*}}\log f^{\alpha,\beta}_{*}*_{v_{*}}M_{\beta})\Big{|}^{2}$
[…]
    $\quad+|v\times(\nabla_{v_{*}}\log f^{\alpha,\beta}_{*}*_{v_{*}}M_{\beta})|^{2}+|v_{*}\times(\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta})|^{2}\Big{)}$
[…]
@@
    $4\frac{|v\times\nabla_{v}f^{\alpha}|^{2}}{(f^{\alpha})^{2}}+4\frac{|v_{*}\times\nabla_{v_{*}}f_{*}^{\alpha}|^{2}}{(f_{*}^{\alpha})^{2}}+|v|^{2}\frac{|\nabla_{v_{*}}f^{\alpha}_{*}|^{2}}{(f^{\alpha}_{*})^{2}}+|v_{*}|^{2}\frac{|\nabla_{v}f^{\alpha}|^{2}}{(f^{\alpha})^{2}},$ (3.39)
@@
    $\lim_{\beta\to 0}\int_{0}^{T}\int_{\mathbb{R}^{12}}\kappa f^{\alpha}f^{\alpha}_{*}|v-v_{*}|^{\gamma}|v\times(\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta})|^{2}{\,\rm d}\eta{\,\rm d}t$ (3.40)
    $=\int_{0}^{T}\int_{\mathbb{R}^{12}}\kappa f^{\alpha}_{*}|v-v_{*}|^{\gamma}\frac{|v\times\nabla_{v}f^{\alpha}|^{2}}{f^{\alpha}}{\,\rm d}\eta{\,\rm d}t$

### v2

@@ 3.2.2 Approximated chain rule
$(\partial_{t}+v\cdot\nabla_{x})\mu^{\alpha,\beta,\delta,\theta}_{t}+\mathsf{E}^{\alpha,\beta,\delta,\theta}_{t}+\frac{1}{2}(\widetilde{\nabla}\cdot\mathcal{U}^{\alpha,\delta,\theta})^{\beta}=0,$ (3.30)
where $(\widetilde{\nabla}\cdot\mathcal{U}^{\alpha,\delta})^{\beta}=(\widetilde{\nabla}\cdot\mathcal{U}^{\alpha,\delta}_{t})*_{(v,v_{*})}M_{\beta}$. The error term $\mathsf{E}^{\alpha,\beta,\delta,\theta}_{t}$ is given by
$\mathsf{E}^{\alpha,\beta,\delta}_{t}=(1-\theta)\int_{\mathbb{R}^{d}}M_{\beta}(u)u\cdot\nabla_{x}f^{\alpha,\delta}_{t}(x,v-u){\,\rm d}u.$
Indeed, we verify (3.30) in the weak formulation with test function $\phi\in C^{\infty}_{c}(\mathbb{R}^{2d})$:
$\frac{d}{dt}\int_{\mathbb{R}^{2d}}\phi{\,\rm d}\mu_{t}^{\alpha,\beta,\delta,\theta}=(1-\theta)\frac{d}{dt}\int_{\mathbb{R}^{2d}}\phi^{\alpha,\beta}{\,\rm d}\mu^{\delta}_{t}+\theta\int_{\mathbb{R}^{2d}}\phi\partial_{t}g{\,\rm d}x{\,\rm d}v,$
@@
$-\int_{\mathbb{R}^{2d}}v\cdot\nabla_{x}\phi{\,\rm d}\mu^{\alpha,\beta,\delta,\theta}_{t}$
$={}$ $-(1-\theta)\int_{\mathbb{R}^{2d}}(v\cdot\nabla_{x}\phi*_{x}M_{\alpha})*_{v}M_{\beta}{\,\rm d}\mu^{\delta}_{t}+\theta\int_{\mathbb{R}^{2d}}\phi(v\cdot\nabla_{x}g){\,\rm d}x{\,\rm d}v$
$={}$ $-(1-\theta)\int_{\mathbb{R}^{2d}}(v\cdot\nabla_{x}\phi*_{x}M_{\alpha}*_{v}M_{\beta}){\,\rm d}\mu^{\delta}_{t}+\theta\int_{\mathbb{R}^{2d}}\phi(v\cdot\nabla_{x}g){\,\rm d}x{\,\rm d}v$
@@
$={}$ $-(1-\theta)\int_{\mathbb{R}^{2d}}(v\cdot\nabla_{x}\phi*_{x}M_{\alpha}*_{v}M_{\beta}){\,\rm d}\mu^{\delta}_{t}+\theta\int_{\mathbb{R}^{2d}}\phi(v\cdot\nabla_{x}g){\,\rm d}x{\,\rm d}v$
$\quad+(1-\theta)\int_{\mathbb{R}^{3d}}\phi(v)M_{\beta}(u)u\cdot\nabla_{x}f^{\alpha,\delta}_{t}(x,v-u){\,\rm d}u{\,\rm d}x{\,\rm d}v,$
@@
$\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\phi^{\alpha,\beta}{\,\rm d}\mathcal{U}^{\delta}_{t}$ $=\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\phi^{\beta}{\,\rm d}\mathcal{U}^{\alpha,\delta}_{t}=-\int_{\mathbb{R}^{2d}}\phi(\widetilde{\nabla}\cdot\mathcal{U}^{\alpha,\delta}_{t})*_{(v,v_{*})}M_{\beta},$
where we use $\Pi_{(v-v_{*})^{\perp}},M_{\alpha}*_{(x,x_{*})}]=0$ and $\cdot$$\cdot$] denotes commutator.
@@
$\leq{}$ $\frac{1}{h}\max\{|\log f^{\alpha,\beta,\delta,\theta}_{t+h}|,|\log f^{\alpha,\beta,\delta,\theta}_{t}|\}|f^{\alpha,\beta,\delta,\theta}_{t+h}-f^{\alpha,\beta,\delta,\theta}_{t}|$
$\leq{}$ $C(\theta)\big(\langle x\rangle+\langle v\rangle\big)\int_{\mathbb{R}}f^{\alpha,\beta}_{r}\Big|\frac{\eta^{\delta}(t+h-r)-\eta^{\delta}(t-r)}{h}\Big|{\,\rm d}r$
$\leq{}$ $C\big(\langle x\rangle+\langle v\rangle\big)\|\eta^{\delta}\|_{\operatorname{Lip}(\mathbb{R})}\int_{0}^{T}f^{\alpha,\beta}_{t}{\,\rm d}t,$
where the bounds on $|\log f^{\alpha,\beta,\delta,\theta}|$ are given in (3.16). The right-hand side of the above inequality is integrable on $\mathbb{R}^{2d}$ by (3.13). We use the dominated convergence theorem to exchange the time derivative and the integral over $\mathbb{R}^{2d}$ to derive
$\frac{d}{dr}\int_{\mathbb{R}^{2d}}f^{\alpha,\beta,\delta,\theta}_{r}\log f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v$
$={}$ $\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta,\delta,\theta}_{r}\partial_{r}f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v+\int_{\mathbb{R}^{2d}}\partial_{r}f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v$
$={}$ $\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta,\delta,\theta}_{r}\partial_{r}f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v.$
@@
$\frac{d}{dt}\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{t})$ $=-\int_{\mathbb{R}^{2d}}\mathsf{E}^{\alpha,\beta,\delta,\theta}_{t}\log f^{\alpha,\beta,\delta,\theta}_{t}{\,\rm d}x{\,\rm d}v$ (3.31)
$\quad+\frac{1}{2}\int_{\mathbb{R}^{4d}}U^{\alpha,\delta,\theta}_{t}\cdot\widetilde{\nabla}(\log f^{\alpha,\beta,\delta,\theta}_{t})^{\beta}{\,\rm d}\eta.$
since $\int_{\mathbb{R}^{2d}}v\cdot\nabla_{x}f^{\alpha,\beta,\delta,\theta}_{t}\log f^{\alpha,\beta,\delta,\theta}_{t}{\,\rm d}x{\,\rm d}v=0$.
@@
$\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{t})-\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{s})=$ $-\int_{s}^{t}\int_{\mathbb{R}^{2d}}\mathsf{E}^{\alpha,\beta,\delta,\theta}_{r}\log f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v{\,\rm d}r$ (3.32)
$+\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}^{4d}}U^{\alpha,\delta,\theta}_{r}\cdot\widetilde{\nabla}(\log f^{\alpha,\beta,\delta,\theta}_{r})^{\beta}{\,\rm d}\eta{\,\rm d}r$
@@
$f^{\alpha,\beta,\delta,\theta}_{t}\to f^{\alpha,\beta,\theta}_{t}\quad\text{uniformly on }(0,T)\times\mathbb{R}^{2d}.$
@@
where $\Omega\subset\mathbb{R}^{2d}$ bounded.
@@
$={}$ $\Big|\int_{\mathbb{R}}f^{\alpha,\beta}_{t-s}\eta^{\delta}(s){\,\rm d}s-\int_{\mathbb{R}}f^{\alpha,\beta}_{t}\eta^{\delta}(s){\,\rm d}s\Big|$
$\leq{}$ $\sup_{s\in\operatorname{supp}(\eta^{\delta})}\Big|\int_{\mathbb{R}^{2d}}(f_{t-s}-f_{t})(y,w)M_{\alpha}(x-y)M_{\beta}(v-w){\,\rm d}y{\,\rm d}w\Big|\to 0$
@@
$\leq{}$ $(\theta g)^{-1}\big|f^{\alpha,\beta,\delta,\theta}_{t}-f^{\alpha,\beta,\theta}_{t}\big|$
$\leq{}$ $C(T,\theta)\big(\langle x\rangle+\langle v\rangle\big)^{a}\big|f^{\alpha,\beta,\delta,\theta}_{r}(x,v)-f^{\alpha,\beta,\theta}_{r}(x,v)\big|,$
@@
$|\nabla_{v}\log f^{\alpha,\beta,\delta}_{t}|\lesssim\beta^{-1}\quad\text{and}\quad|\widetilde{\nabla}(\log f^{\alpha,\beta,\delta}_{t})^{\beta}|\lesssim_{\beta}|v-v_{*}|^{1+\frac{\gamma}{2}}.$ (3.33)
@@
$|\widetilde{\nabla}(\log f^{\alpha,\beta,\delta}_{t})^{\beta}|\lesssim_{\beta,\gamma}|v|^{1+\frac{\gamma}{2}}+|v_{*}|^{1+\frac{\gamma}{2}}.$
@@
The proof follows [CDDW24, Lemma 31 & 32]. We note that
@@
$|\nabla_{v}\log f^{\alpha,\beta,\delta}_{t}|$ $=\Big|\frac{\int_{\mathbb{R}^{d}}f^{\alpha,\delta}(x,w)\nabla_{v}M_{\beta}(v-w){\,\rm d}w}{f^{\alpha,\beta,\delta}_{t}}\Big|$
$\lesssim\beta^{-1}\Big|\frac{\int_{\mathbb{R}^{d}}f^{\alpha,\delta}(x,w)M_{\beta}(v-w){\,\rm d}w}{f^{\alpha,\beta,\delta}_{t}}\Big|\lesssim\beta^{-1}.$
@@
$\widetilde{\nabla}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}=$ $|v-v_{*}|^{1+\frac{\gamma}{2}}\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}-\nabla_{v_{*}}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}\big).$
@@
    $\lim_{\theta\to 0}\lim_{\delta\to 0}\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f_{r}^{\alpha,\beta,\delta,\theta})^{\beta}\cdot U_{r}^{\alpha,\delta,\theta}{\,\rm d}\eta{\,\rm d}r=\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f_{r}^{\alpha,\beta})^{\beta}\cdot U_{r}^{\alpha}{\,\rm d}\eta{\,\rm d}r.$
@@
    $(1-\theta)\Big|\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f_{r}^{\alpha,\beta,\theta})^{\beta}\cdot\Big(U_{r}^{\alpha,\delta}-U_{r}^{\alpha}\Big){\,\rm d}\eta{\,\rm d}r\Big|\to 0\quad\text{and}$ (3.35)
    $(1-\theta)\Big|\int_{s}^{t}\int_{\mathbb{R}^{4d}}\Big(\widetilde{\nabla}(\log f_{r}^{\alpha,\beta,\delta,\theta})^{\beta}-\widetilde{\nabla}(\log f_{r}^{\alpha,\beta,\theta})^{\beta}\Big)\cdot U_{r}^{\alpha,\delta}{\,\rm d}\eta{\,\rm d}r\Big|\to 0.$ (3.36)
@@
    $|\widetilde{\nabla}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}|\lesssim_{\beta}|v-v_{*}|^{1+\frac{\gamma}{2}},\quad|U^{\alpha,\beta,\delta}||v-v_{*}|^{1+\frac{\gamma}{2}}\in L^{1}([0,T]\times\mathbb{R}^{4d}).$ (3.37)
@@
    We define the ball $B^{m}_{k}(\cdot)\subset\mathbb{R}^{md}$ as the ball centred at $\cdot\in\mathbb{R}^{md}$ with radius $k$.
@@
    $|\widetilde{\nabla}(\log f_{r}^{\alpha,\beta,\delta,\theta})^{\beta}-\widetilde{\nabla}(\log f_{r}^{\alpha,\beta,\theta})^{\beta}|$
@@
    $={}$ $\nabla_{v}\Big(\big(\log f^{\alpha,\beta,\delta,\theta}-\log f^{\alpha,\beta,\theta}\big)*_{v}M_{\beta}\Big)$
    $={}$ $\int_{B^{1}_{2l}(0)}\big(\log f^{\alpha,\beta,\delta,\theta}-\log f^{\alpha,\beta,\theta}\big)(u)\nabla_{v}M_{\beta}(v-u){\,\rm d}u$
@@
    $\quad\times\||v-v_{*}|^{1+\frac{\gamma}{2}}U^{\alpha,\delta}\|_{L^{1}([0,T]\times\mathbb{R}^{4d})}\to 0\quad\text{as}\quad\delta\to 0,$
@@
    We use the dominated convergence Theorem 3.13 to show the $\theta$-limit. Notice that the bounds (3.37) imply that $\big|\widetilde{\nabla}\big(\log f^{\alpha,\beta,\theta}\big)^{\beta}\cdot U^{\alpha,\theta}\big|\lesssim\beta^{-1}|U^{\alpha}|$.
@@
    $\lim_{\beta\to 0}\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f^{\alpha,\beta})^{\beta}\cdot U^{\alpha}{\,\rm d}\eta{\,\rm d}r=\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f^{\alpha})\cdot U^{\alpha}{\,\rm d}\eta{\,\rm d}r.$ (3.38)
    We follow [CDDW24]. By Cauchy–Schwarz inequality, we have
    $\Big|\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f^{\alpha,\beta})^{\beta}\cdot U^{\alpha}{\,\rm d}\eta{\,\rm d}r\Big|$
    $\leq\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa f^{\alpha}f^{\alpha}_{*}|\widetilde{\nabla}(\log f^{\alpha,\beta})^{\beta}|^{2}{\,\rm d}\eta{\,\rm d}t+\frac{1}{2}\int_{0}^{T}\mathcal{A}(f^{\alpha}){\,\rm d}t,$
@@
    Similar to (2.16), the product identity $|x|^{2}(y\cdot\Pi_{x^{\perp}}y)=\frac{1}{2}\sum_{i,j=1}^{d}|x_{i}y_{j}-x_{j}y_{i}|^{2}$ implies that
    $\Big|\widetilde{\nabla}(\log f^{\alpha,\beta})^{\beta}\Big|^{2}$
    $=$ ${}\frac{1}{2}\sum_{i,j=1}^{d}|v-v_{*}|^{\gamma}\Big|\Big[v-v_{*},\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}-\nabla_{v_{*}}\log f^{\alpha,\beta}_{*}*_{v_{*}}M_{\beta}\Big]_{ij}\Big|^{2}.$
@@
    $\Big|\Big[v-v_{*},\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}-\nabla_{v_{*}}\log f^{\alpha,\beta}_{*}*_{v_{*}}M_{\beta}\Big]_{ij}\Big|^{2}$
    $\leq{}$ $4\Big(|[v,\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}]_{ij}|^{2}+|[v_{*},\nabla_{v_{*}}\log f_{*}^{\alpha,\beta}*_{v_{*}}M_{\beta}]_{ij}|^{2}$
    $\quad+|[v,\nabla_{v_{*}}\log f^{\alpha,\beta}_{*}*_{v_{*}}M_{\beta}]_{ij}|^{2}+|[v_{*},\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}]_{ij}|^{2}\Big)$
    $\lesssim{}$ $\Big(|[v,\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}]_{ij}|^{2}+|[v_{*},\nabla_{v_{*}}\log f_{*}^{\alpha,\beta}*_{v_{*}}M_{\beta}]_{ij}|^{2}$
@@
    $\frac{|[v,\nabla_{v}f^{\alpha}]_{ij}|^{2}}{(f^{\alpha})^{2}}+\frac{|[v_{*},\nabla_{v_{*}}f_{*}^{\alpha}]_{ij}|^{2}}{(f_{*}^{\alpha})^{2}}+|v|^{2}\frac{|\nabla_{v_{*}}f^{\alpha}_{*}|^{2}}{(f^{\alpha}_{*})^{2}}+|v_{*}|^{2}\frac{|\nabla_{v}f^{\alpha}|^{2}}{(f^{\alpha})^{2}},$ (3.39)
@@
    $\lim_{\beta\to 0}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa f^{\alpha}f^{\alpha}_{*}|v-v_{*}|^{\gamma}|[v,\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}]_{ij}|^{2}{\,\rm d}\eta{\,\rm d}t$ (3.40)
    $=\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa f^{\alpha}_{*}|v-v_{*}|^{\gamma}\frac{|[v,\nabla_{v}f^{\alpha}]_{ij}|^{2}}{f^{\alpha}}{\,\rm d}\eta{\,\rm d}t$